OpenStudy (anonymous):

Simplify square root of 8 divided by fifth root of 8

OpenStudy (anonymous):

\[\sqrt{8}/ \sqrt[5]{8}\]

OpenStudy (anonymous):

is diss stanton ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

are you stanton?

OpenStudy (anonymous):

I'm sorry but I don't know what you mean.

OpenStudy (anonymous):

oh wait no sorry I just re-read it

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

hey can you please help me on this problem i'm really stuck.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Let me think

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

it goes 8 x 8 x 8=512

OpenStudy (anonymous):

there it it

OpenStudy (anonymous):

umm do you want the way only one can come out to be? like the answers?

OpenStudy (anonymous):

8 to the power of negative 7 over 10 8 to the power of negative 1 over 4 8 to the power of 3 over 10 8 to the power of 7 over 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me see somthing

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Wait what grade are you in?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

kk one sec

OpenStudy (anonymous):

It still says 512 in my math :/

OpenStudy (anonymous):

okay thanks can you help me on another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

hat is the exact value for the expression the square root of 171. - the square root of 19. + the square root of 76.? Simplify if possible. 8 the square root of 57. the square root of 19. 2 the square root of 57. 4 the square root of 19.

OpenStudy (anonymous):

*What

OpenStudy (anonymous):

lol one sec

OpenStudy (anonymous):

k sorry :s

OpenStudy (anonymous):

I had to use calculator for this one lol it's 5000211

OpenStudy (anonymous):

What u think?

OpenStudy (anonymous):

im getting confused bc it looks like this. Hang on

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I teach a harvard

OpenStudy (anonymous):

and it's still tough lol

OpenStudy (anonymous):

\[\sqrt{171}-\sqrt{19}+\sqrt{76}\] simplify.

OpenStudy (anonymous):

228

OpenStudy (anonymous):

\[8\sqrt{57} \sqrt{19} 2\sqrt{57} 4\sqrt{19}\]

OpenStudy (anonymous):

its supposed to be one of those.

OpenStudy (anonymous):

lol how much more lol jk

OpenStudy (anonymous):

let's see ok you try and help with this one so you know it really good if it comes up a again you know it ok

OpenStudy (anonymous):

forget it I got it. Thanks though :)

OpenStudy (anonymous):

What was it?

OpenStudy (anonymous):

np